OpenStudy (anonymous):

Please help! I don't understand this problem at all! I need to solve for x :(

OpenStudy (anonymous):

\[\frac{ 2 }{ x-1 }+1 = \frac{ 2 }{ x^2-x }\]

ganeshie8 (ganeshie8):

First notice that \(\large x^2 - x~ =~ x(x-1)\)

ganeshie8 (ganeshie8):

get rid of denominators by multiplying \(\large x(x-1)\) through out : \[\frac{ 2 }{ x-1 }\color{Red}{x(x-1)}+1\color{Red}{x(x-1)} = \frac{ 2 }{ x^2-x }\color{Red}{x(x-1)}\]

ganeshie8 (ganeshie8):

cancelling common factors you get \[\large 2x +x(x-1) = 2\]

ganeshie8 (ganeshie8):

thats just a quadratic equation which can be solved by factoring

OpenStudy (anonymous):

Thanks so much! That helped a lot.

ganeshie8 (ganeshie8):

you're able to solve \(x\) ?